the differential equation of circles passing through the origin and having the centres on the x axis is

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Equation of circle is (x-a)*(x-a) + y^2 = a^2

differentiate once to eliminate constant a

Regards
Vikas TU
4 years ago
Dear student 
x^2 = y^2 +2xydy/dx 
Genera equation of all such circles is 
x^2 + y^2 +2gx = 0 
2x +2ydy/dx +2g = 0 
Desired equation is 
x^2 + y^2 +(-2x-2ydy/dx)x = 0 
So y^2 = x^2 +2xydy/dx 
Hope this helps 
Good Luck